OpenStudy (anonymous):

integral of cotx csc^3x dx

OpenStudy (anonymous):

\[\int\limits \cot x \csc ^3x~dx=\int\limits \csc ^2x \left( \csc x \cot x \right)dx\] \[\csc x=t,~-\csc x~ \cot x~dx=dt\] complete it.